It took a crew 1 h 4 min to row 1 km upstream and back again. If the rate of flow of the stream was 7 km/h, what was the rowing speed of the crew in still water?
Answer by richwmiller(17219)

x+y=1.0666667,
x*(r+7)=1,
y*(r-7)=1
r=8